# Two Minds Better Than One: Collaborative Reward Modeling for LLM Alignment

### v1

@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has made significant progress in aligning large language models (LLMs) with human values Bai:2022:HH-RLHF ; Ouyang:2022:RLHF . One of the core stages in RLHF is reward modeling zhong:2025:RMSurvey ; xu:2025:unifiedseed ; chen:2025:PRM , where reward models (RMs) learn human preferences and intentions based on a pre-collected preference dataset. The insights gleaned from human preferences enable RMs to generate nuanced reward signals, which guide the optimization of the policy model, serving as a proxy during the reinforcement learning phase. Thus, the efficacy of the reward model are crucial for steering LLM to be helpful and harmless Eschmann:2021:reward_design ; pan:2022:reward_effects ; openai:2024:gpt4 .
Pairwise preferences serve as the fuel of reward modeling Wang:2024:secretsII ; liu:2024:SkyworkRewardBagTricks ; zhou:2025:rmb , powering the RM with human intentions and psychological tendencies. However, noisy preferences inevitably arise due to inconsistencies among crowd-sourcing employees. Anthropic and NVIDIA report annotator consistency rates of merely 60%-70% Bai:2022:HH-RLHF ; Wang:2024:helpsteer2 , indicating that human annotators are significantly susceptible to cognitive biases and decision uncertainty.
@@
Recent studies Zheng:2023:MT-Bench ; gao:2024:ImpactPreferenceNoise ; shen:2024:ImprovingReinforcementLearning demonstrate that 20%-40% of open-source preference data is corrupted by noise, with many annotated pairs contradicting the actual human preference. Representative examples selected from HH-RLHF are provided in Appendix A.1. Fig. 1 illustrates that training RMs on noisy preferences significantly degrades downstream RLHF performance and impairs training stability. More details are available in Appendix A.2. This problem can be attributed to reward misgeneralization di:2022:goal_mis ; gao:2023:ScalingLawsReward ; chen:2024:DiscriminRM ; Qiu:2024:RewardTopological , in which distorted signals deviate the LLM away from human-aligned behavior.
Recent work Mitchell:2023:cDPO ; Chowdhury:2024:rDPO ; Wu:2024:DrDPO ; Liang:2024:ROPO attempts to mitigate the impact of noisy preferences on LLM alignment by proposing robust preference optimization objectives. However, these efforts primarily focus on improving training objectives, while neglecting the intrinsic characteristics of preference data. To bridge this gap, we take a data-centric perspective and analyze how noisy preferences affect reward modeling and how their training dynamics differ from clean preferences. Our analysis in Sec. 3.1 reveals that noisy preferences are associated with high training loss and low prediction accuracy. These distinctive attributes suggest the feasibility of identifying and filtering out noisy preferences.
@@ 2 Preliminary
Reward Modeling with RLHF Reward modeling aims to learn human preferences from open-ended conversations. Specifically, the reward model is trained on pre-collected response pairs, where human preferences are explicitly annotated. Following the Bradley-Terry model bradley:1952:BTModel , the probability of preferred response being chosen over the rejected one can be expressed through reward function $r_{\phi}$ as follows:
@@
$\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};\phi\right)=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y_{c}})\sim\mathcal{D}_{\mathrm{rm}}}\left[\log\sigma\left(r_{\phi}\left(\mathbf{y}_{w};\mathbf{x}\right)-r_{\phi}\left(\mathbf{y}_{c};\mathbf{x}\right)\right)\right],$ (2)
@@
where policy model $\pi_{\theta}$ and reference model $\pi_{\mathrm{ref}}$ are initialized from SFT stage, $\beta$ is the coefficient of KL divergence to prevent the $\pi_{\theta}$ deviate significantly from $\pi_{\mathrm{ref}}$, otherwise results in model collapse to single high-reward answers, i.e., reward overoptimization gao:2023:ScalingLawsReward ; Miao:2024:InfoRM .
Direct Preference Optimization DPO Rafailov:2023:DPO directly updates the policy model using offline preference data instead of reward model, deriving a close-form solution for Eq. 3:
@@ 3.1 Motivation
Settings. Experimental analysis is carried out on the HH-RLHF Bai:2022:HH-RLHF preference dataset, which contains approximately 17,000 general multi-round dialogue samples. We follow the standard training pipeline to train RM from scratch and employ the Adam optimizer to minimize the Eq. 2 across $T$ epochs, with the training set randomly shuffled in each epoch. In addition, several metrics are considered to characterize the robustness of preference instances. In particular, we calculate the average loss for individual instance across $T$ epochs, defined as $\mu=\frac{1}{T}\sum_{t=1}^{T}\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};\boldsymbol{\phi}_{t}\right)$. A lower loss indicates strong capability to capture the nuances of preferences and better consistency with annotated preferences, while a higher loss suggests the opposite. We also consider preference accuracy as a more intuitive metric, $acc=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left(r_{\phi_{t}}\left(\mathbf{y}_{w};\mathbf{x}\right)>r_{\phi_{t}}\left(\mathbf{y}_{l};\mathbf{x}\right)\right)$. On the other hand, the variance of loss provides insight into the stability of the training phase, $\sigma=\sqrt{\frac{\sum_{t=1}^{T}\left(\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};\boldsymbol{\phi}_{t}\right)-\mu\right)^{2}}{T}}$, where high variance signifies an ungeneralized pattern, making it challenging for RM to achieve stable reward modeling.
@@ 3.2 Method
Through our empirical analysis, we observe that robust preferences benefit the generalization of the reward model. The use of low-loss as an effective indicator for identifying robust preferences motivates us to enhance reward modeling from a data-centric perspective. Nevertheless, using the self-loss as a filtering criterion will inevitably introduce confirmation bias Palminteri:2017:ConfirmationBias ; Arazo:2020:ConfirmationBias inherited from self-training, verified in the ablation experiments of Sec. 4.3.
@@
$\mathcal{B}_{\psi}=\operatorname{argmax}_{\mathcal{B}_{\phi}\subseteq\mathcal{B},|\mathcal{B}_{\phi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\psi};\mathcal{B}\right),$ (6)
$\mathcal{B}_{\phi}=\operatorname{argmax}_{\mathcal{B}_{\psi}\subseteq\mathcal{B},|\mathcal{B}_{\psi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\phi};\mathcal{B}\right),$ (7)
@@
Extend to the Direct Preference Optimization. Our proposed CRM can seamlessly extend to prevalent implicit-reward alignment methods. By integrating the implicit reward signals from policy LLM, we explore whether policy LLM could achieve improved alignment within our framework when supervisory signal is provided by peer policy. Though our primary analysis focuses on the explicit reward model, we extend our framework on DPO Rafailov:2023:DPO for implicit-reward alignment.
@@ 4.1 Setup
Reward Model Setting. Our experiments utilize Llama3-3B as the base model, and we report preference accuracy on HH-RLHF Bai:2022:HH-RLHF , Ultrafeedback-Binaralized Tunstall:2023:Ultrafeedback and Skywork-Reward liu:2024:SkyworkRewardBagTricks , more details about the dataset can be found in Appendix B.1. Both in-distribution (ID) and out-of-distribution (OOD) test sets are used for evaluation, with RewardBench lambert:2024:rewardbench serving as the OOD test set. Following Liang:2024:ROPO ; gao:2024:ImpactPreferenceNoise , we introduce noise into the training data by randomly flipping response pairs with probabilities of 20% and 40%. In practice, noisy preferences can result from annotation disagreements or malicious attacks.
RLHF Setting. We compare policy LLMs optimized by implicit RM and explicit RM, respectively. The policy LLM is initialized by supervised fine-tuning (SFT) on RLFH-Flow dataset Dong:2024:rlhf , which contains high-quality instruction data sourced from ShareGPT chiang:2023:vicuna , Evol-Instruct xu:2023:evol_instruct , etc. For policy LLMs induced by explicit RM, HH-RLHF is used as a preference dataset for training RM and the prompt data for sampling responses in the RL stage. For policy LLMs induced by implicit RM, we extend our approach to DPO. For implementation details, please refer to Appendix B.2.
Baselines. Our baselines include DPO Rafailov:2023:DPO and other SOTA preference alignment methods to alleviate noisy preferences, including rDPOChowdhury:2024:rDPO , cDPO Mitchell:2023:cDPO and ROPO Liang:2024:ROPO . To suit the reward modeling, we derive their explicit reward objectives in Table 4. A detailed description of the baselines is available in Appendix B.3.
Metrics. We adopt the preference accuracy as a measure of the discriminative ability of RM, specifically indicating that RM’s predictive reward for the preferred response $r_{\phi}\left(\mathbf{y}_{w};\mathbf{x}\right)$ is higher than that of the rejected response, $r_{\phi}\left(\mathbf{y}_{l};\mathbf{x}\right)$. In terms of RLHF, we assess the response quality of policy LLMs. Followed by Rafailov:2023:DPO ; Zheng:2023:Secret_I , we calculate the win-rate by comparing their responses with the SFT targets, where pairs of responses are provided to the superior model GPT-4 for quality comparison. The prompt and evaluation approach within our experiments can be found in Appendix C.4.
@@ 4.2 Main Results
Discriminative Ability of Reward Model. Table 1 presents the discriminative ability of reward model in terms of preference accuracy. Our investigation yields the following observations: (1) As noisy preferences intensify, there is a notable decline in the discriminative capability of the standard RM under both ID and OOD. This phenomenon is evident in the Skywork-Reward, a high-quality preference dataset curated through stringent processes, where preference accuracy deteriorates from $83.85\%$ to $60.60\%$. (2) In the 0% flipped setting, i.e, without the deliberately injection of noise, competitive schemes and CRM outperform the standard RM. This points out the presence of noisy preferences in the original dataset and underscores the inefficiency of the standard RM in handling such noise. (3) From the table, we can observe that CRM consistently achieves superior performance in terms of preference accuracy across various noise levels. In some cases of not the best, CRM also achieves comparable performance with SOTA. Under extreme noise conditions, 40%, our method maintains satisfactory performance, achieving a preference accuracy of 68.23%, compared to cDPO-RM’s 61.90%. Moreover, we supplement more experimental results on RMB zhou:2025:rmb and RM-Bench liu:2025:rmbench in Appendix C.1.
@@
Policies Induced by Explicit Reward Models. To evaluate the effectiveness of RM in RL stage, we compare the policy LLMs optimized with different RMs on two common tasks: general dialogue and summarization. As shown in Table 2, the proposed method significantly enhances response quality in both general dialogue and summarization. Although the original HH-RLHF dataset contains approximately 20%-30% noisy preferences Wang:2024:secretsII , CRM effectively filters underlying noise by peer-reviewing, allowing each model to benefit from robust preferences. This design mitigates the effect of noisy preferences, bolsters reward generalization, and thereby encourages safe and helpful behavior.
Policies Induced by Implicit Reward Models. Following Rafailov:2023:DPO ; Liang:2024:ROPO , we extend our framework for implicit-reward alignment. Table 3 presents win-rate comparisons for competitive baselines, including DPO, cDPO, rDPO and ROPO. Notably, under 20% noise level, the response quality of DPO is significantly compromised by noisy preferences. The win-rates of DPO are 77% and 63% on Anthropic-Helpful and Harmless, respectively, which are worse than the 80% and 65% achieved under 0% noise level. In contrast, our method maintains stability regardless of noisy preferences, underscoring the effectiveness of the collaborative paradigm to refine policy LLM alignment. On average, our framework achieves approximately 67.33% win-rate across different settings, while the second-best achieves 61.55%. Additional results on Arena-Hard li:2024:ArenaHard and MT-Bench Zheng:2023:MT-Bench are shown in Table 9, and qualitative examples are provided in Appendix C.5.
@@ 4.3 Ablation Study and Analysis
Necessity of Peer Review and Curriculum Learning. Recall that Peer Review is designed to provide supervisory feedback on preference selections for the peer model, while Curriculum Learning synchronizes the capabilities of both models to prevent excessive disparities that could impede the effectiveness of peer review. To review the effectiveness of Peer Review and Curriculum Learning, we compare the performance of different variants: (a) w/o Peer Review, (b) w/o Curriculum Learning, (c) w/ Self Review. The former two variants aim to verify the proposed mechanism, while the Self Review variant utilizes feedback from itself, rather than the peer model, to filter noisy preferences. This configuration allows examination of the effects caused by cumulative error and confirmation bias Palminteri:2017:ConfirmationBias ; Arazo:2020:ConfirmationBias . In Fig. 5, the variants exhibit competitive performance while CRM achieves the best results, significantly outperforming each variant. This indicates that our design objectives for Peer Review and Curriculum Learning synergize effectively, enhancing generalization.
Weak Model Reviews Strong Model. Weak to strong transitions burns:2024:weak2strong ; zhu:2025:weaktostrong are scalable in our collaborative framework, prompting us to explore the universality of CRM. In our settings, we utilize two different-size RMs as a strong model and a weak model, then implement the proposed collaborative reward modeling. Fig. 5 presents the performance of two models, illustrating that the collaborative framework significantly enhances the performance of both strong and weak models compared to the standard pipeline. On the one hand, the strong model is endowed with abundant knowledge capacity, guiding the training of the weak model, while the feedback signal from the weak model effectively aids in refining the strong model. Furthermore, we discuss the generalization of the two policy LLMs within our framework in Appendix C.3, underscoring the universality of CRM.

### v2

@@ 1 Introduction
Reinforcement learning from human feedback (RLHF) has made significant progress in aligning large language models (LLMs) with human values [1; 2]. One of the core stages in RLHF is reward modeling [3; 4; 5], where reward models (RMs) learn human preferences and intentions based on a pre-collected preference dataset. The insights gleaned from human preferences enable RMs to generate nuanced reward signals, which guide the optimization of the policy model, serving as a proxy during the reinforcement learning phase. Thus, the efficacy of the reward model are crucial for steering LLM to be helpful and harmless [6; 7; 8].
Pairwise preferences serve as the fuel of reward modeling [9; 10; 11], powering the RM with human intentions and psychological tendencies. However, noisy preferences inevitably arise due to inconsistencies among crowd-sourcing employees. Anthropic and NVIDIA report annotator consistency rates of merely 60%-70% [1; 12], indicating that human annotators are significantly susceptible to cognitive biases and decision uncertainty.
@@
Recent studies [13; 14; 15] demonstrate that 20%-40% of open-source preference data is corrupted by noise, with many annotated pairs contradicting the actual human preference. Representative examples selected from HH-RLHF are provided in Appendix A.1. Fig. 1 illustrates that training RMs on noisy preferences significantly degrades downstream RLHF performance and impairs training stability. More details are available in Appendix A.2. This problem can be attributed to reward misgeneralization [16; 17; 18; 19], in which distorted signals deviate the LLM away from human-aligned behavior.
Recent work [20; 21; 22; 23] attempts to mitigate the impact of noisy preferences on LLM alignment by proposing robust preference optimization objectives. However, these efforts primarily focus on improving training objectives, while neglecting the intrinsic characteristics of preference data. To bridge this gap, we take a data-centric perspective and analyze how noisy preferences affect reward modeling and how their training dynamics differ from clean preferences. Our analysis in Sec. 3.1 reveals that noisy preferences are associated with high training loss and low prediction accuracy. These distinctive attributes suggest the feasibility of identifying and filtering out noisy preferences.
@@ 2 Preliminary
Reward Modeling with RLHF Reward modeling aims to learn human preferences from open-ended conversations. Specifically, the reward model is trained on pre-collected response pairs, where human preferences are explicitly annotated. Following the Bradley-Terry model [24], the probability of preferred response being chosen over the rejected one can be expressed through reward function $r_{\phi}$ as follows:
@@
$\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};r_{\phi}\right)=-\mathbb{E}_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\sim\mathcal{D}}\left[\log\sigma\left(r_{\phi}\left(\mathbf{y}_{w};\mathbf{x}\right)-r_{\phi}\left(\mathbf{y}_{l};\mathbf{x}\right)\right)\right],$ (2)
@@
where policy model $\pi_{\theta}$ and reference model $\pi_{\mathrm{ref}}$ are initialized from SFT stage, $\beta$ is the coefficient of KL divergence to prevent the $\pi_{\theta}$ deviate significantly from $\pi_{\mathrm{ref}}$, otherwise results in model collapse to single high-reward answers, i.e., reward overoptimization [17; 25].
Direct Preference Optimization DPO [26] directly updates the policy model using offline preference data instead of reward model, deriving a close-form solution for Eq. 3:
@@ 3.1 Motivation
Settings. Experimental analysis is carried out on the HH-RLHF [1] preference dataset, which contains approximately 17,000 general multi-round dialogue samples. We follow the standard training pipeline to train RM from scratch and employ the Adam optimizer to minimize the Eq. 2 across $T$ epochs, with the training set randomly shuffled in each epoch. In addition, several metrics are considered to characterize the robustness of preference instances. In particular, we calculate the average loss for individual instance across $T$ epochs, defined as $\mu=\frac{1}{T}\sum_{t=1}^{T}\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};r_{\phi_{t}}\right)$. A lower loss indicates strong capability to capture the nuances of preferences and better consistency with annotated preferences, while a higher loss suggests the opposite. We also consider preference accuracy as a more intuitive metric, $acc=\frac{1}{T}\sum_{t=1}^{T}\mathbb{I}\left(r_{\phi_{t}}\left(\mathbf{y}_{w};\mathbf{x}\right)>r_{\phi_{t}}\left(\mathbf{y}_{l};\mathbf{x}\right)\right)$. On the other hand, the variance of loss provides insight into the stability of the training phase, $\sigma=\sqrt{\frac{\sum_{t=1}^{T}\left(\ell_{\mathrm{BT}}\left(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l};r_{\phi_{t}}\right)-\mu\right)^{2}}{T}}$, where high variance signifies an ungeneralized pattern, making it challenging for RM to achieve stable reward modeling.
@@ 3.2 Method
Through our empirical analysis, we observe that robust preferences benefit the generalization of the reward model. The use of low-loss as an effective indicator for identifying robust preferences motivates us to enhance reward modeling from a data-centric perspective. Nevertheless, using the self-loss as a filtering criterion will inevitably introduce confirmation bias [27; 28] inherited from self-training, verified in the ablation experiments of Sec. 4.3.
@@
$\mathcal{B}_{\psi}=\operatorname{argmax}_{\mathcal{B}_{\psi}\subseteq\mathcal{B},|\mathcal{B}_{\psi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\psi};\mathcal{B}\right),$ (6)
$\mathcal{B}_{\phi}=\operatorname{argmax}_{\mathcal{B}_{\phi}\subseteq\mathcal{B},|\mathcal{B}_{\phi}|=\lambda_{t}|\mathcal{B}|}\sum M\left(r_{\phi};\mathcal{B}\right),$ (7)
@@
Extend to the Direct Preference Optimization. Our proposed CRM can seamlessly extend to prevalent implicit-reward alignment methods. By integrating the implicit reward signals from policy LLM, we explore whether policy LLM could achieve improved alignment within our framework when supervisory signal is provided by peer policy. Though our primary analysis focuses on the explicit reward model, we extend our framework on DPO [26] for implicit-reward alignment.
@@ 4.1 Setup
Reward Model Setting. Our experiments utilize Llama3-3B as the base model, and we report preference accuracy on HH-RLHF [1], Ultrafeedback-Binaralized [29] and Skywork-Reward [10], more details about the dataset can be found in Appendix B.1. Both in-distribution (ID) and out-of-distribution (OOD) test sets are used for evaluation, with RewardBench [30] serving as the OOD test set. Following [23; 14], we introduce noise into the training data by randomly flipping response pairs with probabilities of 20% and 40%. In practice, noisy preferences can result from annotation disagreements or malicious attacks.
RLHF Setting. We compare policy LLMs optimized by implicit RM and explicit RM, respectively. The policy LLM is initialized by supervised fine-tuning (SFT) on RLFH-Flow dataset [31], which contains high-quality instruction data sourced from ShareGPT [32], Evol-Instruct [33], etc. For policy LLMs induced by explicit RM, HH-RLHF is used as a preference dataset for training RM and the prompt data for sampling responses in the RL stage. For policy LLMs induced by implicit RM, we extend our approach to DPO. For implementation details, please refer to Appendix B.2.
Baselines. Our baselines include DPO [26] and other SOTA preference alignment methods to alleviate noisy preferences, including rDPO[21], cDPO [20] and ROPO [23]. To suit the reward modeling, we derive their explicit reward objectives in Table 4. A detailed description of the baselines is available in Appendix B.3.
Metrics. We adopt the preference accuracy as a measure of the discriminative ability of RM, specifically indicating that RM’s predictive reward for the preferred response $r_{\phi}\left(\mathbf{y}_{w};\mathbf{x}\right)$ is higher than that of the rejected response, $r_{\phi}\left(\mathbf{y}_{l};\mathbf{x}\right)$. In terms of RLHF, we assess the response quality of policy LLMs. Followed by [26; 34], we calculate the win-rate by comparing their responses with the SFT targets, where pairs of responses are provided to the superior model GPT-4 for quality comparison. The prompt and evaluation approach within our experiments can be found in Appendix C.4.
@@ 4.2 Main Results
Discriminative Ability of Reward Model. Table 1 presents the discriminative ability of reward model in terms of preference accuracy. Our investigation yields the following observations: (1) As noisy preferences intensify, there is a notable decline in the discriminative capability of the standard RM under both ID and OOD. This phenomenon is evident in the Skywork-Reward, a high-quality preference dataset curated through stringent processes, where preference accuracy deteriorates from $83.85\%$ to $60.60\%$. (2) In the 0% flipped setting, i.e, without the deliberately injection of noise, competitive schemes and CRM outperform the standard RM. This points out the presence of noisy preferences in the original dataset and underscores the inefficiency of the standard RM in handling such noise. (3) From the table, we can observe that CRM consistently achieves superior performance in terms of preference accuracy across various noise levels. In some cases of not the best, CRM also achieves comparable performance with SOTA. Under extreme noise conditions, 40%, our method maintains satisfactory performance, achieving a preference accuracy of 68.23%, compared to cDPO-RM’s 61.90%. Moreover, we supplement more experimental results on RMB [11] and RM-Bench [35] in Appendix C.1.
@@
Policies Induced by Explicit Reward Models. To evaluate the effectiveness of RM in RL stage, we compare the policy LLMs optimized with different RMs on two common tasks: general dialogue and summarization. As shown in Table 2, the proposed method significantly enhances response quality in both general dialogue and summarization. Although the original HH-RLHF dataset contains approximately 20%-30% noisy preferences [9], CRM effectively filters underlying noise by peer-reviewing, allowing each model to benefit from robust preferences. This design mitigates the effect of noisy preferences, bolsters reward generalization, and thereby encourages safe and helpful behavior.
Policies Induced by Implicit Reward Models. Following [26; 23], we extend our framework for implicit-reward alignment. Table 3 presents win-rate comparisons for competitive baselines, including DPO, cDPO, rDPO and ROPO. Notably, under 20% noise level, the response quality of DPO is significantly compromised by noisy preferences. The win-rates of DPO are 77% and 63% on Anthropic-Helpful and Harmless, respectively, which are worse than the 80% and 65% achieved under 0% noise level. In contrast, our method maintains stability regardless of noisy preferences, underscoring the effectiveness of the collaborative paradigm to refine policy LLM alignment. On average, our framework achieves approximately 67.33% win-rate across different settings, while the second-best achieves 61.55%. Additional results on Arena-Hard [36] and MT-Bench [13] are shown in Table 5, and qualitative examples are provided in Appendix C.5.
@@ 4.3 Ablation Study and Analysis
Necessity of Peer Review and Curriculum Learning. Recall that Peer Review is designed to provide supervisory feedback on preference selections for the peer model, while Curriculum Learning synchronizes the capabilities of both models to prevent excessive disparities that could impede the effectiveness of peer review. To review the effectiveness of Peer Review and Curriculum Learning, we compare the performance of different variants: (a) w/o Peer Review, (b) w/o Curriculum Learning, (c) w/ Self Review. The former two variants aim to verify the proposed mechanism, while the Self Review variant utilizes feedback from itself, rather than the peer model, to filter noisy preferences. This configuration allows examination of the effects caused by cumulative error and confirmation bias [27; 28]. In Fig. 5, the variants exhibit competitive performance while CRM achieves the best results, significantly outperforming each variant. This indicates that our design objectives for Peer Review and Curriculum Learning synergize effectively, enhancing generalization.
Weak Model Reviews Strong Model. Weak to strong transitions [37; 38] are scalable in our collaborative framework, prompting us to explore the universality of CRM. In our settings, we utilize two different-size RMs as a strong model and a weak model, then implement the proposed collaborative reward modeling. Fig. 5 presents the performance of two models, illustrating that the collaborative framework significantly enhances the performance of both strong and weak models compared to the standard pipeline. On the one hand, the strong model is endowed with abundant knowledge capacity, guiding the training of the weak model, while the feedback signal from the weak model effectively aids in refining the strong model. Furthermore, we discuss the generalization of the two policy LLMs within our framework in Appendix C.3, underscoring the universality of CRM.